# Second Language (Arabic) Acquisition of LLMs via Progressive Vocabulary Expansion

### v1

@@ Abstract
This paper addresses the critical need for democratizing large language models (LLM) in the Arab world, a region that has seen slower progress in developing models comparable to state-of-the-art offerings like GPT-4 or ChatGPT 3.5, due to a predominant focus on mainstream languages (e.g., English and Chinese). One practical objective for an Arabic LLM is to utilize an Arabic-specific vocabulary for the tokenizer that could speed up decoding. However, using a different vocabulary often leads to a degradation of learned knowledge since many words are initially out-of-vocabulary (OOV) when training starts. Inspired by the vocabulary learning during Second Language (Arabic) Acquisition for humans, the released AraLLaMA employs progressive vocabulary expansion, which is implemented by a modified BPE algorithm that progressively extends the Arabic subwords in its dynamic vocabulary during training, thereby balancing the OOV ratio at every stage. The ablation study demonstrated the effectiveness of Progressive Vocabulary Expansion. Moreover, AraLLaMA achieves decent performance comparable to the best Arabic LLMs across a variety of Arabic benchmarks. Models, training data, benchmarks, and codes will be all open-sourced.
@@
Jianqing Zhu1, Huang Huang2, Zhihang Lin4, Juhao Liang3,4, Zhengyang Tang3,4, Khalid Almubarak5, Abdulmohsen Alharthik1, Bang An1, Juncai He1, Xiangbo Wu3, Fei Yu3,4, Junying Chen3,4, Zhuoheng Ma4, Yuhao Du4, He Zhang4, Emad A. Alghamdi5, Lian Zhang2, Ruoyu Sun3,4, Haizhou Li3,4, Benyou Wangthanks: Benyou Wang is the corresponding author.    These authors contributed equally to this work.3,4, Jinchao Xu1 1 King Abdullah University of Science and Technology, Thuwal, Saudi Arabia 2 Shenzhen International Center for Industrial and Applied Mathematics, Shenzhen Research Institute of Big Data 3 Shenzhen Research Institue of Big Data, Shenzhen, China 4 The Chinese University of Hong Kong, Shenzhen, China 5 King Abdulaziz University, Jeddah, Saudi Arabia
@@ 1 Introduction
In the evolving landscape of large language models (LLMs), the predominant focus has been on English and Chinese. This focus has left other linguistic communities, notably the Arab world, with slower progress in developing comparable models. Within the Arab world  [^1] The Arab World comprises a large group of countries, mainly located in Western Asia and Northern Africa., the development of models such as Jais Sengupta et al. (2023) and AceGPT Huang et al. (2023a) marks a significant step forward, yet these models do not rival the capabilities of state-of-the-art models like GPT-4 or even ChatGPT 3.5. In line with the democratization Touvron et al. (2023), our development of Arabic LLMs focuses on language adaptation settings that utilize existing standard LLM architectures (like LLaMA) and well-trained weights, thereby saving computing resources and ensuring compatibility.
@@
The core challenge in language adaption for English-centric LLMs for a second language is about vocabulary expansion Touvron et al. (2023); Cui et al. (2023); Huang et al. (2023b); Zhao et al. (2024). A case in point is AceGPT Huang et al. (2023b), which struggles with slow decoding speeds due to its inability to adapt to the Arabic vocabulary. It decodes Arabic words into sequences of alphabetical letters rather than at a more efficient granularity, such as Arabic subwords. This inefficiency significantly limits its broader applicability, despite its performance being nearly on par with ChatGPT 3.5 in some benchmarks. The primary challenge associated with vocabulary expansion is the risk that abrupt increases can lead to a high incidence of out-of-vocabulary (OOV) words—words or subwords that are not present in the model’s current vocabulary. Such a surge in OOV words can compromise the linguistic knowledge embedded within the core models. Addressing this issue requires a considerable volume of pre-training data to restore and maintain the model’s linguistic capabilities effectively.
@@
The core philosophy behind AraLLaMA is inspired by the process of vocabulary learning in human Second Language Acquisition, emphasizing that individuals typically expand their vocabulary gradually through incremental learning, rather than through instantaneous acquisition. AraLLaMA progressively extends the Arabic subwords in its vocabulary during pre-training, effectively reducing the ratio of OOV words at every stage. AraLLaMA initialized with LLaMA2 13B, not only seamlessly preserves the inherent knowledge embedded in LLaMA2 13B but also facilitates a smooth transfer of knowledge from English to Arabic. Ablation on TinyLLaMA Zhang et al. (2024) demonstrated the effectiveness of the proposed progressive vocabulary expansion, see Section 6.1.
@@
Followed by extensive instruction tuning, AraLLaMA achieves decent performance comparable to the best Arabic LLMs across various Arabic benchmarks. The contributions of this work are three-fold: 1) We introduce Progressive Vocabulary Expansion, utilizing a modified Byte Pair Encoding (BPE) algorithm inspired by human second language acquisition, and demonstrate its effectiveness. 2) We present AraLLaMA, a pioneering open-source Arabic Large Language Model that decodes Arabic texts three times faster than its predecessor Huang et al. (2023b) while delivering superior performance. 3) We provide the community with access to the complete data processing pipeline, pre-training/fine-tuning data, and model weights. AraLLaMA is compatible with the most popular LLM architecture (i.e., LLaMA) and can be seamlessly integrated into most LLM applications.
@@ 2.1 Cognitively-inspired Motivation: Second Language Acquisition for Humans
Figure 1: Second language acquisition for human, an English-speaking Child’s Journey to Arabic Fluency, From Basic Vocabulary to Cultural Proficiency
@@
In learning a second language (L2), learners pass through several developmental stages as they gain proficiency in L2, including the acquisition of phonetics, vocabulary, grammar, and pragmatic use. Of these language skills, vocabulary acquisition is crucial for language learning. Several studies have posited that L2 learners mostly learn new words incidentally Ramos and Dario (2015); Nation (2001). This suggests that an individual might gradually master a word or a set of words in an unconscious manner. This leads to a phenomenon:
@@ 2.2 Problem Definition: Second Language Acquisition for LLMs
Language adaption
The focus on developing large-scale open-source language models for high-resource languages like English and Chinese has unintentionally marginalized low-resource languages, despite there being about 7,000 languages in use globally. The lack of data and computational resources makes it challenging to develop effective models for these languages. A common practice is to enhance existing models by adding specialized data for these underrepresented languages Cui et al. (2023); Huang et al. (2023b); Zhao et al. (2024), a.k.a, language adaption.
Vocabulary expansion in language adaption
As a preliminary study, we identified Arabic tokens from the LLaMA2 vocabulary using regular expressions. It was observed that the LLaMA2 vocabulary only includes the basic characters of the Arabic language, resulting in relatively slow encoding and decoding speeds compared to English. During domain adaption, it is crucial for vocabulary expansion for the second language, since it could significantly speed up decoding speeds as the number of decoded tokens is reduced due to the adapted vocabulary. Furthermore, although augmenting the existing vocabulary with tokens from additional languages, followed by training on corresponding language corpora, appears to be a logical strategy, empirical evidence suggests that the gains from this method are modest. This insight underscores the complexity of enhancing support for low-resource languages within the framework of current large-scale language models.
@@
Is it beneficial to adopt progressive vocabulary learning in language adaption of LLMs?
3 Methodology: Progressive Vocabulary Expansion for Language Adaption
The standard Byte Pair Encoding (BPE) process expands the initial vocabulary by iteratively merging frequent character pairs or sequences from training data into new tokens, until reaching a desired size. Training commences once this process is completed, rendering the vocabulary static. To investigate the posed question, this section introduces Progressive Vocabulary Expansion. This method incrementally incorporates new tokens in a dynamic vocabulary during training, mimicking a human-like paradigm of digesting and then learning during time.
@@ 3 Methodology: Progressive Vocabulary Expansion for Language Adaption
In contrast to BPE algorithm Sennrich et al. (2015) that uses a static vocabulary during LLM training, we propose an Incremental Byte Pair Encoding (I-BPE) method that uses dynamic vocabulary to implement Progressive Vocabulary Expansion, see Algorithm 1. Similar to the BPE process of repeatedly merging the most frequent pairs, gradually adding new tokens and training them equates to introducing new characters or subwords into the vocabulary, thus expanding and updating it. New tokens are continually added to the vocabulary until the vocabulary size is equal to the given number in each stage, and then the model is trained to adapt to the new vocabulary while increasing the proportion of corpus corresponding to newly added tokens. It repeats this expansion and annealing by gradually increasing both the vocabulary size and proportion of the corresponding corpus until the vocabulary is expanded to a preset size. This iterative approach could improve stability during language adaptation and maintain adaptability to existing data. Technically, this approach could substantially reduce the OOV ratio at every step of the training process, thereby enhancing the model’s capability to gradually recognize previously unknown words.
As seen in Figure 2, there exist two distinct strategies for vocabulary expansion: exponential addition of subwords or uniform addition.
@@
  • The uniform expansion involves adding $K$ tokens at each stage. It results in a total number of $(T-1)\times K$ over $T$ stages while the first stage does not add new tokens.
  • The exponential expansion adds new tokens exponentially, mimicking the vocabulary learning mechanism observed in humans. Consistent with the uniform expansion, there is a stage at the beginning where no new tokens are added and then this approach starts with integrating one new token, with the number of tokens introduced in each subsequent stage doubling, following the sequence $\{0,1,2,\cdots,2^{T-2}\}$, until reaching the desired expansion size.
@@
Figure 2: The impact on compression ratios for uniform or exponential vocabulary expansion.
Exponential Expansion vs. Uniform Expansion
We conduct a comparative analysis of the impact of uniform and exponential vocabulary expansion strategies on token count using the same corpus. The encoding process is segmented into 16 distinct stages, with the token count computed at each stage using the correspondingly expanded vocabulary. Figure 2 illustrates the trend in token counts for both vocabulary expansion methods as the number of stages progresses, Table 9 describes the changes in compression ratio and OOV rate at different stages for the exponential and uniform vocabulary expansion methods. It can be observed that the uniform expansion leads to a significant increase in the compression ratio during the initial stages, but it becomes saturated later on. This could introduce training instability at the beginning, as it suddenly encounters a high ratio of new words, potentially exacerbating catastrophic forgetting in large language models.
In contrast, exponential growth facilitates a gradual adjustment in the compression ratio, maintaining a lower OOV ratio with fewer subwords, which offers a more controlled reduction in the compression ratio as vocabulary size increases. Therefore, we opt for the exponential addition of subwords, as it not only stabilizes the training process but also shortens the length of the decoded sequence by threefold, potentially leading to significant speedups during both training and inference. We set the vocabulary size to be 12,800 words, as this number reaches a saturation point in the compression rate, as shown in Figure 2. To effectively manage OOV rates, we implemented 16 ($\log_{2}^{12800}$) stages of vocabulary expansion based on exponential growth by frequency order.
4 Training
We will discuss the details of data engineering in Section 4.1, along with additional training details in Section 4.2.
@@ 4.1 Data Engineering
Our pre-training dataset comprises both Arabic and English corpora. We employ an array of Arabic corpora encompassing multiple categories as delineated in Table 7 (shown in Appendix C). These include a filtered version of Common Crawl, WebText, and Wikipedia1 sourced from Joud and BAAI, all of which were subjected to an additional cleaning process. Moreover, we gather and methodically purify additional corpora, namely Wikipedia2, Books, and Newspapers. The English corpus is sourced from SlimPajama Soboleva et al. (2023) and Proof-Pile-2 Azerbayev et al. (2023).
Incorporating the insights gained from our discussion on token annealing, this study further delves into the pre-training process, showcasing the integral role of the token annealing strategy in shaping our pre-training stages. Our pre-training framework is meticulously segmented into two epochs, with the inaugural epoch deploying the vocabulary annealing algorithm to fine-tune data distribution, as previously delineated. The subsequent epoch advances with training predicated on the refined vocabulary. The process of vocabulary expansion is methodically organized into 16 delineated stages, each uniquely composed of a calibrated mix of data from English, Arabic, mathematical, and coding domains, with the precise ratios detailed in Table 8 (show in Appendix D). A corpus of 30 billion tokens is employed for training across each stage, underscoring the extensive scale of our pre-training efforts.
The strategic design of these stages showcases a deliberate, phased approach towards the integration of new tokens. This facilitates a seamless adaptation of the model to a broad spectrum of data representations, ensuring a comprehensive understanding and engagement with various linguistic and symbolic nuances. By judiciously modulating the data composition at every stage—wherein the percentage of Arabic data steadily increases, reflecting a focused effort to bolster the model’s proficiency with Arabic, simultaneously with a corresponding decrement in the English data percentage—we guarantee the model’s agility and proficiency across a wide linguistic spectrum.
@@
Specifically, we identified 127 critical topics within Arabic culture, science, and engineering as our focus. ALAN decomposes these topics into a structured hierarchy of fields, sub-fields, and individual disciplines. For each discipline, ALAN compiles a comprehensive list of subjects and designs a syllabus with specific knowledge points for each one. Using GPT-4-0613, ALAN has generated 11,430 subjects and 244,812 detailed knowledge points. We provide more concrete examples in Appendix G.
Armed with this extensive collection of subjects and knowledge points, we direct the LLM to create questions and answers related to these knowledge concepts. The syllabus consists of several lectures, each with 2 to 5 knowledge points. To diversify the knowledge base, we combine knowledge points from both the same and different lectures to produce diverse instructions and answers. Additionally, to vary the instruction types, the LLM generates three kinds of questions at random: multiple-choice, open-ended, and coding questions. In total, we’ve generated 733,419 instruction tuning data pieces using GPT-3.5-Turbo.
We also incorporated instruction tuning data from previous AceGPT projects Huang et al. (2023b), including Quora-Arabic, Alpaca-Arabic Taori et al. (2023), Code-Alpaca-Arabic Chaudhary (2023), Evol-Instruct-Arabic Xu et al. (2023), and ShareGPT data.

### v2

@@ Abstract
This paper addresses the critical need for democratizing large language models (LLM) in the Arab world, a region that has seen slower progress in developing models comparable to state-of-the-art offerings like GPT-4 or GPT-3.5, due to a predominant focus on mainstream languages (e.g., English and Chinese). One practical objective for Arabic LLMs is to utilize Arabic-specific vocabulary in the tokenizer to accelerate decoding. However, using a different vocabulary often leads to degradation of the model’s learned knowledge, since many words become out-of-vocabulary (OOV) at the beginning of training. Inspired by the vocabulary learning during Second Language (Arabic) Acquisition for humans, the released AraLLaMA employs progressive vocabulary expansion, which is implemented by a modified BPE algorithm that progressively extends the Arabic subwords in its dynamic vocabulary during training, thereby balancing the OOV ratio at every stage. The ablation study demonstrated the effectiveness of Progressive Vocabulary Expansion. Moreover, AraLLaMA achieves decent performance comparable to the best Arabic LLMs across a variety of Arabic benchmarks. Our model weights are available at: https://github.com/FreedomIntelligence/AraLLaMa.
@@
Jianqing Zhu1, Huang Huang2, Zhihang Lin4, Juhao Liang3,4, Zhengyang Tang3,4 Khalid Almubarak6, Abdulmohsen Alharthik1, Bang An1, Juncai He1, Xiangbo Wu3, Fei Yu3,4, Junying Chen3,4, Zhuoheng Ma4, Yuhao Du4, He Zhang4, Saied Alshahrani7, Emad A. Alghamdi5, Lian Zhang2, Ruoyu Sun3,4, Haizhou Li3,4, Benyou Wangthanks: corresponding author: wangbenyou@cuhk.edu.cn.    These authors contributed equally to this work.3,4, Jinchao Xu1 1 King Abdullah University of Science and Technology 2 Shenzhen International Center for Industrial and Applied Mathematics, Shenzhen Research Institute of Big Data 3 Shenzhen Research Institute of Big Data 4 The Chinese University of Hong Kong 5 King Abdulaziz University 6 Prince Sattam bin Abdulaziz University 7 University of Bisha
@@ 1 Introduction
In the evolving landscape of large language models (LLMs), the predominant focus has been on English and Chinese. This focus has left other linguistic communities, notably the Arab world, with slower progress in developing comparable models. Within the Arab world  [^1] The Arab World comprises a large group of countries, mainly located in Western Asia and Northern Africa., the development
of models such as Jais Sengupta et al. (2023) and AceGPT Huang et al. (2024) marks a significant step forward, yet these models do not rival the capabilities of state-of-the-art models like GPT-4 Achiam et al. (2023) or even GPT-3.5. In line with democratization Touvron et al. (2023a, b), our development of Arabic LLMs focuses on language adaptation settings that utilize existing standard LLM architectures (like LLaMA Touvron et al. (2023b)) and well-trained weights, thus saving computing resources and ensuring compatibility.
@@
A primary challenge in adapting English-centric LLMs to other languages lies in vocabulary expansion Touvron et al. (2023b); Cui et al. (2023); Huang et al. (2024); Zhao et al. (2024). For instance, AceGPT exhibits slower decoding speeds when processing Arabic, which may be attributed to limitations in its vocabulary adaptation mechanisms. It decodes Arabic words into sequences of alphabetical letters rather than at a more efficient granularity, such as Arabic subwords. This inefficiency significantly limits its broader applicability, despite its performance being nearly on par with GPT-3.5 in some benchmarks. A key concern related to vocabulary expansion is the risk that abrupt increases may result in a high incidence of out-of-vocabulary (OOV) tokens—units absent from the model’s established vocabulary. Such a surge in OOV words can compromise the linguistic knowledge embedded within the core models. Addressing this issue requires a considerable volume of pre-training data to restore and maintain the model’s linguistic capabilities effectively.
@@
The core philosophy behind AraLLaMA is inspired by the process of vocabulary learning in human Second Language Acquisition, emphasizing that individuals typically expand their vocabulary gradually through incremental learning, rather than through instantaneous acquisition. AraLLaMA progressively extends the Arabic subwords in its vocabulary during pre-training, effectively reducing the ratio of OOV words at every stage. AraLLaMA, based on the initialization of LLaMA2 Touvron et al. (2023b), not only retains the foundational knowledge of LLaMA2, but also enables effective cross-lingual transfer from English to Arabic. Ablation on TinyLLaMA Zhang et al. (2024) demonstrated the effectiveness of the proposed progressive vocabulary expansion, see Section 6.1.
@@
Followed by extensive instruction tuning, AraLLaMA achieves decent performance comparable to the best Arabic LLMs across various Arabic benchmarks. The contributions of this work are three-fold: 1) We introduce progressive vocabulary expansion, utilizing a modified byte pair encoding (BPE) algorithm inspired by human Second Language Acquisition, and demonstrate its effectiveness. 2) We present AraLLaMA, a pioneering open-source Arabic Large Language Model that decodes Arabic texts three times faster than its predecessor Huang et al. (2024) while delivering superior performance. 3) We provide the community with access to the complete data processing pipeline, pre-training/fine-tuning data, and model weights. AraLLaMA is compatible with the most popular LLM architecture (i.e., LLaMA) and can be seamlessly integrated into most LLM applications.
@@ 2.1 Cognitively-inspired Motivation: Second Language Acquisition for Humans
Figure 1: Second Language Acquisition for human, an English-speaking child’s journey to Arabic fluency, from basic vocabulary to cultural roficiency
@@
In learning a second language (L2), learners pass through several developmental stages as they gain proficiency in L2, including the acquisition of phonetics, vocabulary, grammar, and pragmatics. Of these language skills, vocabulary acquisition is crucial for language learning. Several studies have posited that L2 learners mostly learn new words incidentally Ramos and Dario (2015); Nation (2001). This suggests that an individual might gradually master a word or a set of words in an unconscious manner. This leads to a phenomenon:
@@ 2.2 Problem Definition: Second Language Acquisition for LLMs
Language adaptation
The focus on developing large-scale open-source language models for high-resource languages like English and Chinese has unintentionally marginalized low-resource languages, despite there being about 7,000 languages in use globally. The lack of data and computational resources makes it challenging to develop effective models for these languages. A common practice is to enhance existing models by adding specialized data for these underrepresented languages Cui et al. (2023); Huang et al. (2024); Zhao et al. (2024), a.k.a, language adaptation.
Vocabulary expansion in language adaptation
As a preliminary study, we identified Arabic tokens from LLaMA2 vocabulary using regular expressions. It was observed that LLaMA2 vocabulary only includes the basic characters of the Arabic language, resulting in relatively slow encoding and decoding speeds compared to English. During domain adaptation, it is crucial for vocabulary expansion for the second language, since it could significantly speed up decoding speeds as the number of decoded tokens is reduced due to the adapted vocabulary. Furthermore, although augmenting the existing vocabulary with tokens from additional languages, followed by training on corresponding language corpora, appears to be a logical strategy, empirical evidence suggests that the gains from this method are modest. This insight underscores the complexity of enhancing support for low-resource languages within the framework of current large-scale language models.
@@
Is it beneficial to adopt progressive vocabulary learning in language adaptation of LLMs?
3 Methodology: Progressive Vocabulary Expansion for Language Adaptation
Conventional Byte Pair Encoding (BPE) algorithms first create a complete vocabulary by iteratively merging the most frequent character pairs from a corpus, and then commence model training with this static vocabulary. This approach, while effective for monolingual models, presents challenges when adapting to new languages as it offers no mechanism for vocabulary evolution during training. To address this limitation, we propose Progressive Vocabulary Expansion.
3.1 Incremental Byte Pair Encoding Algorithm
In contrast to standard BPE algorithms Sennrich et al. (2015) that use a static vocabulary, we introduce Incremental Byte Pair Encoding (I-BPE) that dynamically augments the vocabulary during training. This approach mirrors human language acquisition, where vocabulary growth occurs simultaneously with deepening language comprehension. Algorithm 1 outlines our method.
@@ 3.1 Incremental Byte Pair Encoding Algorithm
The key innovation of I-BPE is its staged approach to vocabulary expansion. At each stage, we expand the vocabulary to a predetermined size, then train the model while gradually increasing the proportion of data corresponding to newly added tokens. This approach significantly reduces out-of-vocabulary (OOV) tokens at each training phase, enabling the model to incorporate new linguistic elements while preserving previously acquired knowledge.
3.2 Expansion Strategies Comparison
For implementing vocabulary expansion, we investigated two principled strategies (illustrated in Figure 2):
@@ 3.2 Expansion Strategies Comparison
Uniform Expansion: Adds a fixed number $K$ of tokens at each stage, resulting in $(T-1)\times K$ total additions over $T$ stages.
Exponential Expansion: Doubles the number of new tokens at each stage following the sequence $\{0,1,2,\cdots,2^{T-2}\}$, mimicking human language acquisition patterns.
@@
Figure 2: Compression ratio comparison between uniform and exponential vocabulary expansion strategies.
Our comparative analysis using an identical Arabic corpus through 16 progressive stages revealed crucial differences between the two approaches. As shown in Figure 2 and detailed in Table 11, uniform expansion causes abrupt improvements in compression ratio during early stages followed by diminishing returns. This sudden introduction of many tokens creates training instabilities and risks catastrophic forgetting as the model’s representation space must rapidly accommodate numerous new tokens simultaneously.
Exponential expansion, however, offers critical advantages through its graduated approach: it provides superior training stability as the gradual introduction of tokens allows smooth adaptation of the model’s representation space; it maintains consistent OOV ratios throughout training, preventing sudden vocabulary distribution shifts; and it achieves significant computational efficiency with a 3 times reduction in sequence length compared to the original LLaMA tokenizer. Based on these findings, we implemented exponential expansion with 12,800 Arabic subwords across 16 stages ($\log_{2}^{12800}$), representing the optimal saturation point for compression ratio improvement.
3.3 Compression Ratios and Tokenizer Evaluation
To rigorously assess the effectiveness of our vocabulary expansion approach, we conducted a comprehensive comparative evaluation of tokenization performance across multiple leading models. Using an identical Arabic corpus of 39 million words, we analyzed how different tokenizers processed Arabic text, with particular attention to efficiency metrics that impact both performance and computational requirements.
Tokenizer Total Words Total Tokens Subword Fertility Ratio of Words Unbroken Rényi Efficiency
LLaMA2(AceGPT) 39,006,442 210,027,671 5.3844 0.0183 0.7731
Bloomz 39,006,442 80,617,499 2.0668 0.3176 0.7709
Mistral 39,006,442 206,082,344 5.2833 0.0185 0.7928
Jais 39,006,442 75,126,494 1.9260 0.3895 0.7343
Our model 39,006,442 66,554,771 1.7063 0.6323 0.7491
Table 1: Comprehensive tokenizer evaluation using standard metrics across different models.
The results in Table 1 reveal notable differences in how these models handle Arabic text. Our model achieved a token compression ratio of 0.3174 (ratio of tokens to original text size), representing a 68% improvement over LLaMA2’s baseline, which directly enhances inference speed and reduces memory requirements. We evaluated several key metrics established in recent tokenizer evaluation literature:
  1. 1.
    Subword Fertility Rust et al. (2021); Moosa et al. (2023): This metric measures the average number of tokens per word. Our model achieves the most optimal fertility (1.7063), approximately 3 times more efficient than LLaMA2 (5.3844) and Mistral (5.2833), while also outperforming multilingual models like Bloomz Muennighoff et al. (2022) (2.0668) and Jais (1.9260) that were specifically designed with Arabic support.
  2. 2.
    Word Integrity Moosa et al. (2023): For Arabic’s rich morphology, preserving words as single tokens is vital. Our model achieves 63.23% word integrity, far exceeding LLaMA2 (1.8%) and outperforming Arabic-optimized models like Jais (38.95%) and Bloomz (31.76%).
  3. 3.
    Total Tokens: For the identical test corpus, our model requires only 66.55 million tokens, compared to LLaMA2’s 210.03 million,a reduction of approximately 68% that translates directly to memory savings and computational efficiency in both training and inference phases.
  4. 4.
    Rényi Efficiency Zouhar et al. (2023): This information-theoretic measure (higher values indicate better vocabulary utilization) shows our model (0.7491) achieves comparable efficiency to LLaMA2 (0.7731) despite its much lower token count, indicating efficient use of vocabulary space while maintaining high word integrity.
The comparative analysis indicates that our model achieves an optimal equilibrium between morphological preservation and computational efficiency. Although models such as LLaMA2 and Mistral exhibit marginally superior Rényi Efficiency coefficients, this advantage is offset by substantial deficiencies in word integrity preservation and significantly elevated token densities. When compared with models specifically optimized for Arabic processing, such as Jais and Bloomz, our model consistently demonstrates superior performance across the majority of evaluation metrics, validating the efficacy of the progressive vocabulary expansion methodology for non-Latin script languages.
4 Training Methodology
Based on the Progressive Vocabulary Expansion methodology described above, we develop AraLLaMA, an Arabic Large Language Model that implements our proposed I-BPE algorithm. In this section, we detail the AraLLaMA training process, including data engineering and training specifics.
@@ 4.1 Data Engineering
Our pre-training dataset comprises both Arabic and English corpora. We employ an array of Arabic corpora encompassing multiple categories as delineated in Table 9 (shown in Appendix D). These include filtered versions of Common Crawl, WebText, and Wikipedia1 sourced from Joud and BAAI, all of which were subjected to additional cleaning processes. Moreover, we gather and purify additional corpora, namely Wikipedia2, Books, and Newspapers. The English corpus is sourced from SlimPajama Soboleva et al. (2023) and Proof-Pile-2 Azerbayev et al. (2023).
@@
Specifically, we identified 127 critical topics within Arabic culture, science, and engineering as our focus. ALAN decomposes these topics into a structured hierarchy of fields, sub-fields, and individual disciplines. For each discipline, ALAN compiles a comprehensive list of subjects and designs a syllabus with specific knowledge points for each one. Using GPT-4-0613, ALAN has generated 11,430 subjects and 244,812 detailed knowledge points. We provide more concrete examples in Appendix H.
Armed with this extensive collection of subjects and knowledge points, we direct the LLM to create questions and answers related to these knowledge concepts. The syllabus consists of several lectures, each with 2 to 5 knowledge points. To diversify the knowledge base, we combine knowledge points from both the same and different lectures to produce diverse instructions and answers. Additionally, to vary the instruction types, the LLM generates three kinds of questions at random: multiple-choice, open-ended, and coding questions. In total, we’ve generated 733,419 instruction tuning data pieces.
We also incorporated instruction tuning data from previous AceGPT projects. These include Quora-Arabic, Alpaca-Arabic Taori et al. (2023), Code-Alpaca-Arabic Chaudhary (2023), Evol-Instruct-Arabic Xu et al. (2023), and ShareGPT data.